OpenStudy (anonymous):

Find the area A and circumference C of a circle of radius 3ft. Express the answer in terms of pi

OpenStudy (anonymous):

A=pi*r^2=pi*3^2=9*pi sq ft. C=2*r*pi=6*pi feet